A random variable follows the continuous uniform distribution between 15 and 35. ​a) Calculate the probabilities below for the distribution. ​1) ​P(x≤30​) ​2) ​P(x=33​) ​b) What are the mean and standard deviation of this​ distribution?
Answer #1

99f xn Gorica,b) » PIX <x) ; aca26 :. P( x < 30) - 3 = 15/ao = 0.| 6 P (x=33) = 10 (Prob. at any point =0 because no area © m